# Resolving Availability and Run-time Integrity Conflicts in Real-Time Embedded Systems

### v1

@@ Abstract.
Run-time integrity enforcement in real-time systems presents a fundamental conflict with availability. Existing approaches in real-time systems primarily focus on minimizing the execution-time overhead of monitoring. After a violation is detected, prior works face a trade-off: (1) prioritize availability and allow a compromised system to continue to ensure applications meet their deadlines, or (2) prioritize security by generating a fault to abort all execution.
In this work, we propose PAIR, an approach that offers a middle ground between the stark extremes of this trade-off. PAIR monitors real-time tasks for run-time integrity violations and maintains an Availability Region (AR) of all tasks that are safe to continue. When a task causes a violation, PAIR triggers a non-maskable interrupt to kill the task and continue executing a non-violating task within AR. Thus, PAIR ensures only violating tasks are prevented from execution, while granting availability to remaining tasks. With its hardware approach, PAIR does not cause any run-time overhead to the executing tasks, integrates with real-time operating systems (RTOSs), and is affordable to low-end microcontroller units (MCUs) by incurring +2.3% overhead in memory and hardware usage. [^1] This work is currently under peer review.
@@ 1. Introduction
Embedded systems have become ubiquitous, often serving important roles in critical infrastructure responsible for real-time sensing or actuation, where deterministic and timely responses are crucial for functionality and safety. They are implemented with microcontroller units (MCUs), low-powered resource-constrained processing units optimized for energy efficiency. MCUs typically lack advanced features, such as memory management units (MMUs) or the ability to run rich operating systems (OS) that facilitate inter-process isolation. To make up for this, real-time operating systems (RTOSs) can be incorporated as an additional software layer to provide basic task management (e.g., scheduling, inter-task communication, timing services) with low overhead. Still, limited architectural protections make MCUs vulnerable (nafees2023smart).
@@
Methods like secure boot (lohr2010patterns) ensure the code integrity at boot-time, but additional measures are required to provide integrity guarantees at run-time. For example, an adversary ($\sf{\mathcal{A}dv}$) could inject code into memory, then exploit a vulnerability (e.g., a buffer overflow (cowan2000buffer)) to redirect execution to the injected code. Although these code-injection-based control flow attacks can be mitigated with memory protection techniques (e.g., $W\oplus X$ and DEP (ms_dep)), they do not protect against code-reuse attacks like return- or jump-oriented programming (ROP/JOP) (rop; jop), that chain together carefully selected instruction gadgets from unmodified code. To account for such attacks, run-time integrity mechanisms are required.
A prominent run-time integrity mechanism is Control Flow Integrity (CFI) (cfi-survey-1), which employs reference monitors to detect when control flow instructions (e.g., calls, returns, jumps) deviate from an expected control flow path. Typically, valid paths are determined by constructing a Control Flow Graph (CFG) via static analysis and comparing the destination of a control flow transition with the set of valid locations in the CFG. Since the exact destination of some branch instructions (e.g., returns, indirect jumps/calls) cannot be determined from static analysis, mechanisms like shadow stacks (burow2019sok), type-checking (type_check_cfi), and landing pads or pointer authentication (arm_manual_bti_pac) can be used to narrow the scope of valid destinations.
@@
Recent CFI methods for embedded and real-time systems (cfi_rts_survey) enable cost-effective enforcement of valid control flows. Several designs (ecfi; recfish; insectacide; pac-pl; hcfi; eilid) focus on performing CFI checks in a way that minimizes run-time overhead on a program (or task) while monitoring integrity. However, most works overlook the handling of tasks after a CFI violation has been raised. On the other hand, availability-focused security techniques do not account for run-time violations (garota; de2022casu), delay intervention until task completion (ecfi), or invoke the scheduler without task integrity guarantees (aion). Given this trade-off, it is unclear from current works how to achieve a middle ground between post-violation integrity and availability.
@@
To bridge this gap, we introduce PAIR: Preserving Availability and Integrity at Run-time. PAIR demonstrates that both availability and run-time integrity can be prioritized simultaneously. PAIR hardware components monitor run-time integrity of independent tasks within a given program. Then, PAIR dynamically adds and removes them from an Availability Region (AR) based on their integrity status. When any run-time integrity violation is detected, PAIR aborts execution of the violating task, removes it from AR, and continues executing another task in AR. PAIR does not allow a task to return to AR until a trusted software update occurs. To summarize, we claim the following contributions:
@@
    formally defined properties of PAIR security, its modules sub-properties, and demonstration via model-checking that the sup-properties and security goals are upheld (Sec. 4);
@@
    an open-source FPGA prototype of PAIR (repo), evaluated with a real-world RTOS and example benchmark applications, demonstrating applicability to low-end MCUs due to minimal memory and hardware overheads (Sec. 7).
@@ 2. Background
2.1. Scope
We consider embedded systems operating real-time or sensing-based tasks atop low-end, single-core MCUs, with limited addressable memory (e.g., 8-256KB), split into program memory (PMEM) and data memory (DMEM). They are typically 8- or 16-bit CPUs running at 1-16MHz (e.g., TI MSP430). Such devices usually run software at bare metal, meaning they lack MMUs, and thus perform memory accesses and execute instructions from their physical addresses. We assume an RTOS is implemented for task scheduling and timing services.
@@ 2.2. Control Flow Integrity (CFI)
Many techniques employ their reference monitor through software instrumentation (bin-cfi; ccfir; lockdown; typearmor), enabling deployment of CFI without hardware changes. Alternatively, hardware extensions that enable CFI, such as Branch Target Identification (BTI) and Pointer Authentication (PA) in ARM (arm_manual_bti_pac) and Intel Control-Flow Enforcement Technology (CET) (intel-cet) in commercial processors, have led to subsequent proposals for fully fledged CFI schemes (pac-ret; pacstack; fineibt; pal). ARM has made BTI and PA available among their class of embedded devices, along with hardware tracing extensions that have also been leveraged to deploy CFI in embedded systems (sherloc; insectacide).
2.3. Real-Time Operations in Embedded Systems
Embedded systems can implement an RTOS that efficiently provides critical services for real-time operations (e.g., task scheduling, inter-task communication, timing adherence). After an application assigns tasks priority and schedules them through an RTOS API, the RTOS manages their execution. To do so, the RTOS contains functions to control execution, such as killing a task or yielding to another task. A task runs until it voluntarily yields or until a time period has elapsed, at which point the RTOS intervenes to perform a context switch and resume the next highest-priority task.
Many works have proposed security mechanisms that account for real-time operations in embedded systems, including mechanisms to provide authenticated (iscflat; pearts; rap-track; asap) and audited (acfa; traces) proofs of valid execution, availability mechanisms for trusted software (garota; aion; de2022casu), and local run-time integrity in the form of CFI (cfi_rts_survey). CFI for embedded systems is either fully software-based (ecfi; recfish; pac-pl) or uses hardware extensions with minimal software support (eilid; sherloc; insectacide). Regardless of the approach, all aim to minimize disruptions to the executing task during the CFI monitoring. Upon the violation, they abort execution to prioritize run-time integrity, compromising the availability of a system in an unknown state. However, in the case of real-time systems, aborting execution system-wide can be costly. An alternative approach (ecfi) is to allow the system to log the violation and continue executing. Although this prioritizes availability, it could be costly to allow a system in an unknown state to continue execution.
@@ 2.4. Linear Temporal Logic
We use Linear Temporal Logic (LTL) (ltl) to specify a formal model of PAIR hardware and the properties that it should satisfy. In Sec. 7, we use a model checking (specifically NuSMV (nusmv)) to determine whether the system adheres to the specified properties. Similarly to prior works (ucca; vrased; garota; apex; asap), we describe hardware using LTL specifications containing propositional connectives (e.g., conjunctions $\land$, disjunctions $\lor$, negation $\neg$, implication $\rightarrow$) and temporal connectives (e.g., $X\phi$ to denote a condition $\phi$ is true at the next system state , and $G\phi$ to denote a condition $\phi$ globally holds for all future states) between system states.
@@ 3. PAIR at a High Level
[P1] Post-Violation Availability: when a task has a run-time integrity violation, PAIR always interferes to (1) remove the task from AR and (2) yield to another non-violating task;
@@
A high-level overview of an PAIR-enabled system is depicted in Fig. 1. It assumes a system that operates real-time tasks (denoted $T_{i}$) with an RTOS, while PAIR introduces two trusted software components. Initially, the MCU runs with no violations, and PAIR has assigned all tasks to AR. PAIR monitors MCU signals and memory to determine which real-time task is executing at any given moment. PAIR interfaces with an IM to detect the violation when it occurs (1). Upon this detection, PAIR will generate a non-maskable interrupt (2) to invoke the kill-and-yield trampoline, while also removing the violating task from AR. These steps are repeated for any other tasks that generate a violation. The bounds of AR are not expanded again until a trusted software update has completed (4), and at this point PAIR will reinstate all violating tasks back into AR.
@@ 4.1. Adversary Model
We consider an adversary ($\sf{\mathcal{A}dv}$) that exploits software vulnerabilities to manipulate any software or data, unless it is explicitly protected by hardware. $\sf{\mathcal{A}dv}$ aims to cause run-time integrity violations (e.g., control flow hijack and ROP/JOP attacks (jop; rop)) to invoke malicious actions or avoid critical actions. $\sf{\mathcal{A}dv}$ aims to continue its malicious actions even after detection, hoping to prevent intervention or other tasks from executing. Hardware attacks that require physical attacks to circumvent hardware protections are out of scope for this work, as they require orthogonal physical access control measures (obermaier2018past).
@@ 4.2. Hardware and Memory Configuration
    $D_{addr}$: the memory address that is targeted by the current instruction (i.e., the address being written to/read from if $W_{en}$$R_{en}$);
@@
Figure 2. System Configuration: PAIR interfaces with the MCU to read specific signals pertaining to the currently executing instruction (e.g., PCPC, $W_{en}$, $R_{en}$, $D_{addr}$, irqirq).
@@
    $\mathsf{D_{\textsf{{PAIR}}}}$: contains all critical data referenced by SW or hardware.
@@ 5.1. PAIR module Sub-properties
Task Memory Access Integrity: (2) $G:((R_{en}\lor W_{en})\land(D_{addr}\in D_{\textsf{{PAIR}}})\land\neg(PC\in SW))\rightarrow violation_{\textsf{{PAIR}}})$ (3) $G:(W_{en}\land(D_{addr}\in PMEM)\land\neg(PC\in SW)\rightarrow violation_{\textsf{{PAIR}}})$ (4) $G:(violation_{\textsf{{PAIR}}}\rightarrow X(trigger))$
@@
LTL 5 and 6 specify the requirements to control PAIR’s non-maskable interrupt signal trigger. LTL 5 shows that trigger should be set in the next state when $violation_{IM}$ has been set by the IM. LTL 6 shows that trigger is cleared in the next state when trigger is set in the current and the irq has been set, denoting that the interrupt has been accepted by the MCU.
@@ 5.2. PAIR goals via sub-properties
Fig. 7 outlines the theorems that state properties [P1] and [P2] are upheld. These theorems are also checked via model checker, described further in Sec. 7. Here we also discuss the intuition behind why these theorems uphold the security properties.
@@
Theorem 1 states that LTLs 2, 3, 4, 5, and 6 uphold Definition 1 for [P1]. Definition 1 specifies that each instance of $violation_{IM}$ or $violation_{\textsf{{PAIR}}}$ must always result in trigger in the next state. It is provided by LTLs 5 and 6 that trigger is set at this instance and is not cleared until the MCU generates its interrupt signal, denoting it was accepted. Additionally, LTLs 2 and 3 ensure that a task’s memory accesses that are relevant to PAIR result in trigger, regardless of the specific IM in use.

### v2

@@ Abstract.
Run-time integrity enforcement in real-time systems presents a fundamental conflict with availability. Existing approaches in real-time systems primarily focus on minimizing the run-time overhead of monitoring. After a violation is detected, prior works face a trade-off: (1) prioritize availability and allow a compromised system to continue to ensure applications meet their deadlines, or (2) prioritize security by generating a fault to abort all execution.
In this work, we propose PAIR, an approach that offers a middle ground between the stark extremes of this trade-off. PAIR monitors real-time tasks for run-time integrity violations and maintains an Availability Region (AR) of all tasks that are safe to continue. When a task causes a violation, PAIR triggers a non-maskable interrupt to kill the task and continue executing a non-violating task within AR. Thus, PAIR ensures only violating tasks are prevented from execution, while granting availability to remaining tasks. With its hardware approach, PAIR does not cause any run-time overhead to the executing tasks. PAIR’s design is formally verified, integrates with real-time operating systems (RTOSs), and is affordable to low-end microcontroller units (MCUs) by incurring +2.3% overhead in memory and hardware usage.
@@ 1. Introduction
Embedded systems have become ubiquitous, often serving important roles in critical infrastructure responsible for real-time sensing or actuation, where deterministic and timely responses are crucial for functionality and safety. They are implemented with microcontroller units (MCUs), low-powered resource-constrained processing units optimized for energy efficiency. MCUs typically lack advanced features, such as memory management units (MMUs) or the ability to run rich operating systems (OS) that facilitate inter-process isolation. To make up for this, real-time operating systems (RTOSs) can be incorporated as an additional software layer to provide basic task management (e.g., scheduling, inter-task communication, timing services) with low overhead. Still, limited architectural protections make MCUs vulnerable (Nafees et al., 2023).
@@
One prevalent threat includes run-time attacks, in which an adversary ($\sf{\mathcal{A}dv}$) could inject code into memory, then exploit a vulnerability (e.g., a buffer overflow (Cowan et al., 2000)) to redirect execution to the injected code. Although memory-protection techniques can mitigate code-injection-based control flow attacks, they do not protect against code-reuse attacks like return- or jump-oriented programming (ROP/JOP) (Roemer et al., 2012; Bletsch et al., 2011), that chain together carefully selected instruction gadgets from unmodified code. To account for such attacks, run-time integrity mechanisms are required.
@@
A prominent run-time integrity mechanism is Control Flow Integrity (CFI) (Burow et al., 2017), which employs reference monitors to detect when control flow instructions (e.g., calls, returns, jumps) deviate from an expected control flow path. Recent CFI methods for embedded and real-time systems (Mishra et al., 2022) enable cost-effective enforcement of valid control flows. Several designs (Tan and Zhao, 2023; Walls et al., 2019; Wang et al., 2024; Serra et al., 2022; Christoulakis et al., 2016; Jakkamsetti et al., 2025) focus on performing CFI checks in a way that minimizes run-time overhead on a program (or task) while monitoring integrity. However, most works overlook the handling of tasks after a CFI violation has been raised. On the other hand, availability-focused security techniques for MCUs (Aliaj et al., 2022; De Oliveira Nunes et al., 2022; Abbasi et al., 2017; Alder et al., 2021; Caulfield et al., 2023, 2024) do not account for run-time violations (Aliaj et al., 2022; De Oliveira Nunes et al., 2022; Alder et al., 2021) or delay intervention until task completion (Abbasi et al., 2017) or verification (Caulfield et al., 2023, 2024; Lengert et al., 2026). Given this trade-off, it is unclear from current works how to achieve a middle ground between post-violation integrity and availability.
@@
To bridge this gap, we introduce PAIR: Preserving Availabi- lity and Integrity at Run-time. PAIR demonstrates that both availability and run-time integrity can be prioritized simultaneously. PAIR hardware components monitor run-time integrity of independent tasks within a given program. Then, PAIR dynamically adds and removes them from an Availability Region (AR) based on their integrity status. When any run-time integrity violation is detected, PAIR aborts execution of the violating task, removes it from AR, and continues executing another task in AR. PAIR does not allow a task to return to AR until a trusted software update occurs. To summarize, we claim the following contributions:
@@
    formally defined properties of PAIR security, its modules sub-properties, and demonstration via model-checking that security goals are upheld (Secs. 5- 6);
@@
    a FPGA prototype of PAIR, evaluated with a real-world RTOS and example benchmark applications, demonstrating applicability to low-end MCUs due to minimal memory and hardware overheads (Sec. 8).
@@ 2.1. Control Flow Integrity (CFI)
Many techniques employ their reference monitor through software instrumentation (Burow et al., 2017; Walls et al., 2019; Jakkamsetti et al., 2025), enabling deployment of CFI without hardware changes. Alternatively, hardware extensions that enable CFI, such as Branch Target Identification (BTI) and Pointer Authentication (PA) in ARM (Arm Ltd., 2025), have led to subsequent proposals for fully fledged CFI schemes (ARM, 2018; Liljestrand et al., 2021). ARM has made BTI and PA available among their class of embedded devices, along with hardware tracing extensions that have also been leveraged to deploy CFI in embedded systems (Tan and Zhao, 2023; Wang et al., 2024).
2.2. Real-Time Operations in MCUs
For real-time operations, MCU software can use an RTOS for critical services (e.g., task scheduling, inter-task communication, timing adherence). After an application assigns tasks priority and schedules them through an RTOS, the RTOS manages their execution. To do so, the RTOS contains functions to control execution, such as killing a task or yielding to another task. Tasks run until they voluntarily yield or until a time period has elapsed, at which point the RTOS performs a context switch and resumes the next highest-priority task.
Many works have proposed security mechanisms that account for real-time operations in embedded systems, including mechanisms to provide authenticated (Neto and Nunes, 2023; Neto et al., 2024; Caulfield et al., 2022) and audited (Caulfield et al., 2023, 2024; Lengert et al., 2026) proofs of valid execution, availability mechanisms for trusted software (Aliaj et al., 2022; Alder et al., 2021; De Oliveira Nunes et al., 2022), and local run-time integrity in the form of CFI (Mishra et al., 2022). CFI for embedded systems is either fully software-based (Abbasi et al., 2017; Walls et al., 2019; Serra et al., 2022) or uses hardware extensions with minimal software support (Jakkamsetti et al., 2025; Tan and Zhao, 2023; Wang et al., 2024). Regardless of the approach, all aim to minimize disruptions to the executing task during the CFI monitoring. Upon the violation, they abort execution to prioritize run-time integrity, compromising the availability of a system in an unknown state. However, in the case of real-time systems, aborting execution system-wide can be costly. An alternative approach from prior works is to log the violation for later handling (Abbasi et al., 2017). However, this is equally unacceptable in this scenario, as it allows the hijacked task to continue with no guarantees regarding safety.
@@ 2.3. Linear Temporal Logic
We use Linear Temporal Logic (LTL) (Vardi, 2005) to specify a formal model of PAIR hardware and the properties that it should satisfy. In Sec. 8, we use a model checking (NuSMV (Cimatti et al., 2002)) to determine whether the system adheres to the specified properties. Similarly to prior works (Aliaj et al., 2022; Caulfield et al., 2022), we describe hardware using LTL specifications containing propositional connectives (e.g., conjunctions $\land$, disjunctions $\lor$, negation $\neg$, implication $\rightarrow$) and temporal connectives (e.g., $X\phi$ to denote a condition $\phi$ is true at the next system state , and $G\phi$ to denote a condition $\phi$ globally holds for all future states) between system states.
2.4. Scope
We consider embedded systems operating real-time or sensing-based tasks atop low-end, single-core MCUs, with limited addressable memory (e.g., 8-256KB), split into program memory (PMEM) and data memory (DMEM). They are typically 8- or 16-bit CPUs running at 1-16MHz (e.g., TI MSP430). Such devices usually run software at bare metal, meaning they lack MMUs, and thus perform memory accesses and execute instructions from physical addresses.
We consider the lowest end of MCUs, on which hardware features for access control (Ltd., 2026; Instruments, 2015) are either limited or non-existent. For the sake of this work, we assume the latter. However, we do assume the MCU has an integrity monitor (e.g., CFI support) as this is widely studied and already deployed into real-world MCUs (Arm Ltd., 2025). We consider implications of more advanced platforms in Section 10.
In terms of software, we assume modularized set of tasks execute on the MCU. Each task has differences in timing requirements and priority. To manage these task characteristics, we assume an RTOS is also in use for scheduling, timing services, and synchronization between the tasks.
3. Application Example
Consider a simple MCU that is used to implement a syringe pump controller that allows physicians to monitor and deliver medication to patients remotely over the network (Islam et al., 2019; Saqlain et al., 2023). In this scenario, the device may be installed with multiple software submodules serving different roles, operating as separate tasks atop an RTOS for prioritization, scheduling, and timely execution. For example, it may include the following:
  • $\bullet$
    a WiFi module for network communication,
  • $\bullet$
    LCD display and buzzer control for user interface and feedback,
  • $\bullet$
    sensor monitoring of patient vitals (e.g., temperature and heart-rate monitoring), and
  • $\bullet$
    motor control for accurate and timely dosage injection.
Suppose the network communication task contains a buffer overflow (Cowan et al., 2000) on the stack, enabling a control flow hijack attack (e.g., ROP (Roemer et al., 2012)). Given a CFI integrity monitor has been installed in the device, attempts to exploit this vulnerability will be detected. As discussed in Section 2, current approaches would not offer a suitable balance: they would either suspending all tasks for the duration of the reset and recovery window, including motor control and sensor monitoring, or allow the compromised network communication module to continue without safety guarantees.
Scenarios like this demand a mechanism that can simultaneously prevent violating tasks while ensuring non-violating tasks can continue their time-sensitive execution. Critically, hardware-based mechanisms are required, as other software (e.g., tasks or the RTOS itself) may attempt to reschedule the violating task before it is properly remediated.
@@ 4. PAIR at a High Level
[P1] Post-Violation Availability: when a task has a run-time integrity violation, PAIR always (1) removes the task from AR and (2) triggers the yielding to a non-violating task;
@@
A high-level overview of an PAIR-enabled system is depicted in Fig. 1. It assumes a system that operates real-time tasks (denoted $T_{i}$) with an RTOS, while PAIR introduces two trusted software components. Initially, the MCU runs with no violations, and PAIR has assigned all tasks to AR. PAIR monitors MCU signals and memory to determine which real-time task is executing at any given moment. PAIR interfaces with an IM to detect the violation when it occurs (1). Upon this detection, PAIR will generate a non-maskable interrupt (2) to invoke the kill-and-yield trampoline, while also removing the violating task from AR. These steps are repeated for any other tasks that generate a violation. The bounds of AR are not expanded until a trusted software update has completed (4), and at this point PAIR will reinstate all violating tasks back into AR.
@@ 5.1. Adversary Model
We consider an adversary ($\sf{\mathcal{A}dv}$) that exploits software vulnerabilities to manipulate any data or software (tasks or the RTOS), unless it is explicitly protected by hardware. $\sf{\mathcal{A}dv}$ aims to cause run-time integrity violations (e.g., control flow hijack and ROP/JOP attacks (Bletsch et al., 2011; Roemer et al., 2012)) to invoke malicious actions or avoid critical actions. $\sf{\mathcal{A}dv}$ aims to continue its malicious actions even after detection by re-entering revoked tasks. Physical attacks to circumvent hardware protections are out of scope for this work, as they require orthogonal physical access control measures (Obermaier and Immler, 2018).
@@ 5.2. Hardware and Memory Configuration
    $D_{addr}$: the memory address targeted by the current instruction (i.e., the address being written/read if $W_{en}$$R_{en}$);
@@
    $\mathsf{D_{\textsf{{PAIR}}}}$: all critical data referenced by SW or hardware.
@@
Figure 2. System Configuration: PAIR interfaces with MCU, an Integrity Monitor (IM), and memory
@@ 6.1. PAIR module Sub-properties
Task Memory Access Integrity: (2) $G:\{(R_{en}\lor$ $W_{en})\land(D_{addr}\in D_{\textsf{{PAIR}}})$ $\land\neg(PC\in SW)\rightarrow violation_{\textsf{{PAIR}}})\}$ (3) $G:\{W_{en}\land$ $(D_{addr}\in PMEM)$ $\land\neg(PC\in SW)\rightarrow violation_{\textsf{{PAIR}}}\}$ (4) $G:(violation_{\textsf{{PAIR}}}\rightarrow X(trigger))$
@@
LTL 5 and 6 specify the requirements to control PAIR’s non-maskable interrupt signal trigger. LTL 5 shows that trigger should be set in the next state when $violation_{IM}$ has been set by the IM. LTL 6 shows that trigger is cleared in the next state when trigger is currently set and the irq has been set, denoting that the interrupt has been accepted by the MCU.
@@ 6.2. PAIR goals via sub-properties
Fig. 7 outlines the theorems that state properties [P1] and [P2] are upheld. These theorems are also checked via model checker, described further in Sec. 8. Here we also discuss the intuition behind why they are upheld.
@@
Theorem 1 states that LTLs 2, 3, 4, 5, and 6 uphold Definition 1 for [P1]. Definition 1 specifies that each instance of $violation_{IM}$ or $violation_{\textsf{{PAIR}}}$ must always result in trigger in the next state. It is provided by LTLs 5 and 6 that trigger is set at this instance and is not cleared until the MCU generates its interrupt signal, denoting it was accepted. Additionally, LTLs 2 and 3 ensure that a task’s memory accesses relevant to PAIR result in trigger, regardless of the specific IM in use.